# Live Fitness Coaching as a Testbed for Situated Interaction

### v1

@@ 4 Baseline Stream-VLM
Current state of the art vision-language models [13, 30, 32, 38, 60, 63] are largely “turn-based” – they take as input a image/video along with an instruction and provide a text output. Our QEVD-FIT-COACH benchmark instead requires models to provide feedbacks proactively, i.e., without explicit prompting, based solely on the actions of the participants in a streaming setting. To this end, we propose a baseline streaming video language model, Stream-VLM, specialized to the fitness coaching domain. It consists of a 3D CNN-based vision backbone [1, 40] for understanding fine-grained fitness actions and a LLaMA-2-based language backbone [53]. Special action tokens are introduced to enable feedback delivery without explicit prompting. We begin by describing the vision backbone in detail, followed by the special action tokens and the training scheme.
Vision backbone. It is designed to robustly recognize motion cues, crucial for fitness coaching. This is in contrast to current state-of-the-art video-language models [38, 60] which typically use CLIP/ViT based vision encoders which capture scene content rather than motion. In detail, our vision backbone architecture is based on a publicly available 3D CNN [1, 40] which was shown to be able to recognize a wide range of behavior patterns, including simple exercises. It consists of a mix of 2D and 3D convolutional layers, ensuring that the model can pick up both motion and content information of individual frames for predictions—both of which are relevant to provide appropriate feedbacks. Additionally, the convolutional layers are causal making it well-suited for the streaming setting of our QEVD-FIT-COACH benchmark. Features from the vision backbone are fused into the LM backbone through cross attention at several layers following the methodology of [2, 9].
@@
Training scheme. Our Stream-VLM streaming baseline is trained end-to-end in three stages: (1) The vision backbone (3D CNN) is pre-trained on ImageNet[16] followed by the QEVD-FIT-300K short-clips video collection described in Sec. 3.2; (2) Next, the model is trained end-to-end on the fitness questions and fitness feedbacks annotations (excluding the long-range videos) from the QEVD-FIT-COACH dataset. The purpose of this stage is to align the vision backbone (3D CNN) and LM with the pre-trained action recognition capability of the vision backbone. Hence, only the adapter (cross-attention layer) weights are updated; (3) Finally, the model is fine-tuned on long-range videos from the QEVD-FIT-COACH fitness feedbacks subset with feedback annotations interleaved to reflect a interactive streaming setting. We limit the model training to $30$ second individual exercise segments and leave it to future work to train across the full workout. Note, the LLaMA-2 backbone is fine-tuned using LoRA (dim = 64) [25]. In pactice, we found it helpful to reduce the weight on the <next> token to $0.1$ (in the cross entropy loss) due to the long input video sequences under consideration. Additional details are provided in the appendix.
@@
InstructBLIP [13] 0.047 0.040 0.839 1.64
Video-LLaVA [32] 0.057 0.025 0.847 1.82
Video-ChatGPT [38] 0.098 0.078 0.850 2.27
Video-LLaMA [60] 0.101 0.077 0.859 2.28
LLaMA-VID [30] 0.100 0.079 0.859 2.33
LLaVA-Next [63] 0.104 0.078 0.858 2.39
@@
Socratic-Llama-2-7B 0.094 0.071 0.860 2.39 0.50
Video-ChatGPT [38] 0.108 0.093 0.863 2.42 0.50
LLaMA-VID [30] 0.106 0.090 0.860 2.40 0.50
Stream-VLM 0.125 0.116 0.863 2.56 0.59
Stream-VLM (w/o 3D CNN) 0.090 0.083 0.857 2.17 0.51
Stream-VLM (w/o Action-Tokens) 0.125 0.110 0.861 2.56 0.50
@@ 5 Experiments
Fluency. We use the METEOR [7], ROUGE-L [33] and BERT [62] metrics to evaluate fluency. For these metrics, we first temporally match predicted and ground truth feedbacks. Each ground truth response is matched to the closest predicted response within a $3$ second window, maintaining their temporal order. The respective METEOR, ROUGE-L and BERT scores are computed on only the matched feedbacks.
Automated evaluation (LLM-Accuracy). In addition to the metrics above, we employ an LM for holistic feedback evaluation [35, 38]. In contrast to the metrics above, LMs offer the advantage that they better correlate with human preferences. We provide the LM with ground truth and predicted feedbacks. The LM then scores the predicted feedbacks holisticallly for accuracy. We use the state-of-the-art open-source Mixtral-Instruct-0.1 [26] LLM, with scores in the range 1-5.
Temporal F-Score. To measure whether predicted responses occur at the correct time-step we compute a temporal F-score. Predicted responses within a session are classified as true or false positives based on whether they temporally match ground truth responses as described for the fluency metrics above. Predicted responses with a ground truth match are true positives, otherwise false positives. Ground truth responses without a match are false negatives. This allows us to calculate temporal precision, recall, and thus the temporal F-Score.
@@ 5.2 Evaluation on the QEVD-FIT-COACH Benchmark
We begin by evaluating state of the art (open source) video-language models InstructBLIP [13], Video-LLaVA [32], Video-ChatGPT [38], Video-LLaMA [60], LLaMA-VID [30], LLaVA-Next[34] on the QEVD-FIT-COACH benchmark (see Tab. 3). As these models are “turn-based” – they cannot respond interactively to an input video – we prompt them to provide feedbacks at regular intervals. Specifically, to remain faithful to the streaming setting in our QEVD-FIT-COACH benchmark, we always prompt these models (zero-shot) with the entire video, including a history of generated feedbacks, up to the latest time-step. We use an interval of $5$ seconds, which is the average silence period within an exercise segment in the QEVD-FIT-COACH dataset. While LLaMA-VID [30] and LLaVA-Next [34] perform best, overall performance across all zero-shot models is weak as shown in Tab. 3. We present qualitative examples in Fig. 4 which highlight repetitive and uninformative feedbacks provided by LLaMA-VID and LLaVA-Next. They are unable to provide corrective feedback at the right time largely due to their lack of fitness domain knowledge and their “turn-based” nature.
@@
Next, we address the lack of fitness domain knowledge by fine-tuning Video-ChatGPT and LLaMA-VID on QEVD following the process discussed in Sec. 4. As shown in Tab. 4, we see that fine-tuning significantly improves performance as expected. However, the performance gain is still limited by the CLIP/ViT based visual encoders which are not well suited for representing fine-grained human motion, on top of being turn-based.
@@
To deal with these issues, the Stream-VLM baseline uses a 3D CNN trained to recognize fine-grained fitness activities and special action tokens to enable interactive feedback. We also consider the following ablations of the Stream-VLM model: (1) instead of the 3D CNN, we use a CLIP based encoder similar to Video-ChatGPT [38] (“w/o 3D CNN”); (2) we use a non-interactive turn-based version without the <next> and <feedback> action tokens (“w/o Action Tokens”). We also consider a text-only Socratic model [59]: Socratic-Llama-2-7B, where we prompt the language-only LLaMA-2-7B LM to generate a feedback for the previous $5$ seconds of user activity provided as a list of activity descriptions in $1$-second intervals. Similar to the zero-shot video-language model evaluations, the full history of described activity and generated feedbacks is included in the prompt. The textual description of user activity is based on the activations of the aforementioned fine-tuned 3D-CNN. An example prompt is provided in the appendix.
@@
The results in Tab. 4 demonstrate that the Stream-VLM model surpasses the performance of the other models. Crucially, we see a significant improvement in the temporal F-score (0.59 vs 0.50) in comparison to the turn-based models. This is also illustrated in Fig. 4, where the Stream-VLM model is able to provide corrective feedbacks at the right time compared to the turn-based baselines. The improved quality of the feedbacks is also reflected in the METEOR, ROUGE-L and LLM-Accuracy metrics. Finally, the advantage of the 3D CNN is illustrated both by the weak performance of the Stream-VLM (w/o 3D CNN) ablation and the strong performance of the Socratic-Llama-2-7B baseline – where an off-the-shelf LLaMA-2-7B model is able to outperform state of the art vision-language models using just the activations of the 3D CNN as a prompt. Overall, while there is significant room for improvement, these results suggest that end-to-end training is a viable path towards good performance on our QEVD-FIT-COACH benchmark and more broadly on the task of responding interactively to events in the visual stream.
Figure 4: Predicted feedbacks on the FIT-Coach benchmark. The “turn-based” LLaMA-VID and LLaVA-Next models are unable to provide corrective feedback and instead provide only very generic and repetitive feedback. The Stream-VLM models has learned to provide appropriate feedback at the appropriate time.
@@ 6 Conclusion
We propose QEVD, a novel interactive visual coaching benchmark and dataset as a test-bed for real-time, real-world situated interaction, and show that this task is challenging for existing LLM-based architectures. As a first step towards closing the gap to situation interaction, we also present Stream-VLM, a streaming vision-language model baseline that learns not only what to say, but also when to say it, based on user activity in the incoming video stream. Overall, we consider our work as a starting point for research into end-to-end training of domain-specific interactive vision models, and hope that our data and baselines will encourage further work in this area.
Privacy and ethics. The data was collected under a direct agreement with the crowd workers, permitting research and commercial use. Furthermore, a detector was used on all videos to detect any issues, such as individuals in the background, followed by manual inspection of the videos that scored above a threshold, to remove such videos from the collection. Personal identifiable information from the videos was removed to the extent possible, e.g., audio and meta-data. All participants received appropriate and fair compensation for the regions where the participants were located.
@@
Broader Impact. In addition to potential bias mentioned in the previous paragraph, language models can produce harmful and biased content, make incorrect claims and produce wrongful advice. This needs to be taken into account when interacting with, deploying or building on these models, particularly in sensitive domains like fitness coaching where incorrect advice may lead to physical harm. While the grounding in visual input supports the generation of language that is contextual, it is not a remedy against these deficiencies of language models. It also has to be taken into account that any computer vision model processing visual information about human subjects could in principle extract information beyond what is required for the use-case, such as biometric information.
@@ Appendix A Overview
Here we provide, 1. Additional qualitative examples from QEVD-FIT-300K and QEVD-FIT-COACH; 2. Additional data collection and annotation details including instructions provided to the crowdworkers; 3. Additional training details of the Stream-VLM model; 4. Details of the prompts used for the (zero-shot) baselines in Tab. 3, including the prompt used to compute the LLM-Accuracy described in Sec. 5.1.
@@ Appendix B Additional Qualitative Examples from QEVD-FIT-300K
In Fig. 5, we provide additional example annotations from the fitness questions and fitness feedbacks on short-clip videos. We see that high-level questions focus on overall aspects of the video, e.g., “What exercise is the user doing? Describe how they are doing it.”. On the other hand, fine-grained questions focus on details such as the user’s rate of punch (Fig. 5 top) or whether the user’s back is rounded (Fig. 5 bottom). The feedbacks on these short-clips provide positive reinforcement, e.g., “This is looking great!” if the user is doing the exercise correctly (Fig. 5 top) and also provide useful advice, e.g., toe alignment while performing squats (Fig. 5 bottom).
Figure 5: Additional example annotations available on the short video clips from the QEVD-FIT-300K dataset (see also Figs. 2 and 1 in the main paper)
@@ Appendix C Additional Data Collection Details
Here we provide additional details pertaining to the various subsets of QEVD.
@@ C.1 QEVD-FIT-300K
Here we discuss the labels associated with the short video clips in the QEVD-FIT-300Kdataset. These labels are crucial for the generation of the fitness questions and feedbacks for for short-clip videos.
@@
Coarse and fine labels. As described in Sec. 3.2 in the main paper, the short video clips are annotated with coarse labels and fine-grained attributes in addition to the questions and feedbacks. The full list of coarse labels and fine-grained attributes can be found at https://developer.qualcomm.com/software/ai-datasets/qevd.
@@ C.2 QEVD-FIT-COACH
Video Collection. Long-range workout videos were collected by letting users perform workout sessions consisting of three $1$-minute sections, where each section consisted of two exercises. Candidate exercises per section are shown in Tab. 5. Users were instructed to perform the exercise shown on screen for the indicated duration while facing the camera to imitate a virtual fitness coaching setup. They were also instructed to perform common mistakes and their corrections (similar to the short-clips). While exercising, users were also presented temporally-smoothed top-k activations (showing mistakes and feedbacks) from the pre-trained vision model (but they were free to ignore this). The recorded videos were subsequently annotated with fully contextual feedback using a simple web interface with careful attention to preserve causal interactions.

### v2

@@ 4 Baseline Stream-VLM
Current state-of-the-art vision-language models [15, 33, 35, 42, 65, 68] are largely turn-based—they take an image or video as input along with an instruction and produce a textual output. In contrast, our Qevd-Fit-Coach benchmark requires models to provide feedback proactively, i.e., without explicit prompting, based solely on the participants’ actions in a streamed setting. To this end, we propose a baseline streaming video language model, Stream-VLM, specialized to the fitness coaching domain. It consists of a 3D-CNN-based vision backbone [1, 44] for understanding fine-grained fitness actions and a LLaMA-2-based language backbone [58]. Special action tokens are introduced to enable feedback delivery without explicit prompting. We begin by describing the vision backbone in detail, followed by the special action tokens, and finally, the training scheme.
Vision backbone. Our vision backbone is designed to robustly recognize motion cues crucial for fitness coaching. This is in contrast to current state-of-the-art video-language models [42, 65], which typically use CLIP-/ViT-based vision encoders to capture scene content rather than motion. Specifically, our architecture is based on a publicly available 3D CNN [1, 44] capable of recognizing a wide range of behavior patterns, including simple exercises. It consists of a mix of 2D and 3D convolutional layers, ensuring that the model can pick up both motion and content information of individual frames to make predictions—both of which are relevant to provide appropriate feedbacks. Additionally, the convolutional layers are causal, making the model well-suited for the streaming setting of our Qevd-Fit-Coach benchmark. Features from the vision backbone are fused into the LM backbone through cross attention at several layers following the methodology of [4, 11].
@@
Training scheme. Our Stream-VLM streaming baseline is trained end-to-end in three stages: (1) The vision backbone (3D CNN) is pre-trained on ImageNet[18], followed by the Qevd-Fit-300K short-clips video collection described in Sec. 3.2; (2) Next, the model is trained end-to-end on the fitness questions and fitness feedbacks annotations (excluding the long-range videos) from the Qevd-Fit-Coach dataset. The purpose of this stage is to align the vision backbone (3D CNN) and LM with the pre-trained action recognition capability of the vision backbone. Hence, only the adapter (cross-attention layer) weights are updated; (3) Finally, the model is fine-tuned on long-range videos from the Qevd-Fit-Coach fitness feedbacks subset with feedback annotations interleaved to reflect an interactive streaming setting. We limit the model training to $30$-second individual exercise segments and leave it to future work to train on workouts spanning multiple exercises. The LLaMA-2 language backbone is fine-tuned using LoRA (dim = 32) [27]. The 3D CNN and adapter (cross-attention layer) weights are kept frozen. Additional details are provided in the appendix.
@@
InstructBLIP [15] 0.047 0.040 0.839 1.56
Video-LLaVA [35] 0.057 0.025 0.847 2.16
Video-ChatGPT [42] 0.098 0.078 0.850 1.91
Video-LLaMA [65] 0.101 0.077 0.859 1.29
LLaMA-VID [33] 0.100 0.079 0.859 2.20
LLaVA-NeXT [68] 0.104 0.078 0.858 2.27
@@
Socratic-LLaMA-2-7B 0.094 0.071 0.860 2.17 0.50
Video-ChatGPT [42] 0.108 0.093 0.863 2.33 0.50
LLaMA-VID [33] 0.106 0.090 0.860 2.30 0.50
Stream-VLM 0.125 0.110 0.864 2.44 0.54
Stream-VLM (w/o 3D CNN) 0.090 0.083 0.857 2.11 0.51
Stream-VLM (w/o Pre-training) 0.095 0.087 0.858 2.08 0.52
Stream-VLM (w/o Action-Tokens) 0.125 0.110 0.861 2.41 0.50
@@ 5 Experiments
Fluency. We use the METEOR [9], ROUGE-L [36] and BERT [67] metrics to evaluate fluency. The METEOR, and ROUGE-L metrics assess lexical similarity between the ground truth and predicted feedbacks: e.g., in the case of a corrective feedback where the person is not moving their arms, these metrics would prefer predicted feedbacks referring to the “arm” and “not moving“. The BERT score on the other hand matches feedbacks at a semantic level.
To compute these metrics, we first temporally match predicted and ground truth feedbacks. Each ground truth response is matched to the closest predicted response within a $3$ second window, maintaining their temporal order. The respective METEOR, ROUGE-L and BERT scores are then computed on only the matched feedbacks.
Automated evaluation (LLM-Accuracy). In addition to the metrics above, we employ an LLM for holistic feedback evaluation [39, 42]. In contrast to the metrics above, LLMs offer the advantage that they better correlate with human preferences [70]. We provide the LM with ground truth and predicted feedbacks. The LM then scores the predicted feedbacks holistically for accuracy. We use the state-of-the-art open-source LLaMA-3-70B-Instruct [3] LLM, with scores in the range $1$ to $5$. This LLM-Accuracy metric along with METEOR, ROUGE-L and BERT metrics ensures that the predicted feedbacks match the ground-truth both at the semantic level while containing references to specific important terms.
Temporal F-Score. To measure temporal accuracy we assess whether predicted responses occur at the correct time-step and compute a temporal F-score. Predicted responses are classified as true or false positives based on whether they temporally match ground truth responses as described for the fluency metrics. Predicted responses without a ground truth match are false positives and ground truth responses without a matching predicted response are false negatives. This allows us to calculate temporal precision, recall, and hence, the temporal F-Score.
Figure 4: Predicted feedbacks on the Fit-Coach benchmark. The “turn-based” LLaMA-VID and LLaVA-NeXT models are unable to provide corrective feedback and instead generate overly generic and repetitive feedback. The Stream-VLM model has learned to provide relevant feedback at the appropriate time.
@@ 5.2 Evaluation on the Qevd-Fit-Coach Benchmark
We begin by evaluating state-of-the-art (open source) video-language models, including InstructBLIP [15], Video-LLaVA [35], Video-ChatGPT [42], Video-LLaMA [65], LLaMA-VID [33], and LLaVA-NeXT[38], on the Qevd-Fit-Coach benchmark (see Tab. 3). Since these models are “turn-based”—they cannot respond interactively to an input video—we prompt them to provide feedbacks at regular intervals. Specifically, to remain faithful to the streaming setting in our Qevd-Fit-Coach benchmark, we always prompt these models (zero-shot) with the entire video, including a history of generated feedbacks, up to the latest time-step. We use an interval of $5$ seconds, equivalent to the average silence period within an exercise segment in the Qevd-Fit-Coach dataset. While LLaMA-VID [33] and LLaVA-NeXT [38] perform best among zero-shot baselines, overall performance across all zero-shot models is weak as shown in Tab. 3. We present qualitative examples in Fig. 4, highlighting the repetitive and uninformative nature of the feedback provided by LLaMA-VID and LLaVA-NeXT. They are unable to provide corrective feedback at the right time largely due to their lack of fitness domain knowledge and their “turn-based” nature.
@@
Next, we address the lack of fitness domain knowledge by fine-tuning Video-ChatGPT and LLaMA-VID on Qevd following the process discussed in Sec. 4. As shown in Tab. 4, fine-tuning significantly improves performance as expected. However, the performance gain is still limited by the CLIP-/ViT-based visual encoders, which are not well-suited for representing fine-grained human motion, not to mention the limitations incurred from their turn-based nature.
@@
To deal with these issues, our Stream-VLM baseline uses a 3D CNN trained to recognize fine-grained fitness activities and special action tokens to enable interactive feedback. We also consider the following ablations of the model: (1) instead of the 3D CNN, we use a CLIP-based encoder similar to Video-ChatGPT [42] (“w/o 3D CNN”); (2) we skip pre-training the Stream-VLM with the Qevd-Fit-300K short-clips fitness questions and feedbacks dataset (“w/o Pre-training”); (3) we use a non-interactive turn-based version without the <next> and <feedback> action tokens (“w/o Action Tokens”). We also consider a text-only Socratic model, Socratic-LLaMA-2-7B [64]. In this model, we prompt the language-only LLaMA-2-7B LLM to generate feedback for the previous $5$ seconds of user activity, provided as a list of activity descriptions in $1$-second intervals. Similar to the zero-shot video-language model evaluations, the full history of described activity and generated feedbacks is included in the prompt. The textual description of user activity is based on the activations of the aforementioned fine-tuned 3D CNN. An example prompt is provided in the appendix.
@@
The results in Tab. 4 demonstrate that the Stream-VLM model surpasses the performance of the other models. Crucially, we see a significant improvement in the temporal F-score (0.59 vs 0.50) in comparison to the turn-based models. This is also illustrated in Fig. 4, where the Stream-VLM model is shown to provide relevant corrective feedback at the appropriate time as opposed to the turn-based baselines. The improved quality of the feedbacks is also reflected in the METEOR, ROUGE-L and LLM-Accuracy metrics. The drop in both fluency and temporal accuracy resulting from the pre-training ablation (w/o Pre-training) supports the quality and utility of the fitness feedbacks and questions within our Qevd-Fit-300K dataset. Furthermore, the advantage of the 3D CNN is demonstrated by two observations: the weak performance of Stream-VLM when the 3D CNN is ablated (w/o 3D CNN), and the strong performance of the Socratic-LLaMA-2-7B baseline. In the latter, an off-the-shelf LLaMA-2-7B model outperforms state-of-the-art vision-language models using only the activations of the 3D CNN as a prompt.
Overall, while there is significant room for improvement, these results suggest that end-to-end training is a viable path towards good performance on our Qevd-Fit-Coach benchmark and more broadly, on the task of responding interactively to events within a visual stream.
@@ 6 Conclusion
We propose Qevd, a novel interactive visual coaching benchmark and dataset, as a test-bed for real-time, real-world situated interaction, and demonstrate that this task is challenging for existing LLM-based architectures. As a first step towards closing the gap to situated interaction, we present Stream-VLM, a streaming vision-language model baseline that learns not only what to say, but also when to say it, based on user activity in the incoming video stream. Overall, we consider our work a starting point for research into end-to-end training of domain-specific interactive vision models, and hope that our data and baselines will encourage further work in this area.
Privacy and ethics. The data was collected under a direct agreement with the crowd workers, permitting research and commercial use. Furthermore, a detector was used on all videos to detect any issues, such as individuals in the background, followed by manual inspection of the videos that scored above a threshold, to remove such videos from the collection. Personal identifiable information from the videos was removed to the extent possible, e.g., audio and meta-data. Participants received appropriate and fair compensation for the regions where they were located.
@@
Broader Impact. In addition to potential bias mentioned in the previous paragraph, language models can produce harmful and biased content, make incorrect claims and produce wrongful advice. This needs to be taken into account when interacting with, deploying or building on these models, particularly in sensitive domains like fitness coaching where incorrect advice may lead to physical harm. Although the grounding in visual input supports the generation of language that is contextual, it is not a remedy against these deficiencies of language models. It is also important to consider that any computer vision model processing visual information about human subjects could, in principle, extract information beyond what is required for the use-case, such as biometric information.
@@ Appendix A Overview
Here we provide: (1) Additional qualitative examples from Qevd-Fit-300K and Qevd-Fit-Coach; (2) Additional data collection and annotation details, including instructions provided to the crowd-workers; (3) Additional training details of the Stream-VLM model; (4) Details of the prompts used for the zero-shot baselines in Tab. 3, including the prompt used to compute the LLM-Accuracy described in Sec. 5.1.
@@ Appendix B Additional Qualitative Examples from Qevd-Fit-300K
In Fig. 5, we provide additional annotation examples from the fitness questions and fitness feedbacks on short-clip videos. As shown, high-level questions focus on overall aspects of the video, e.g., “What exercise is the user doing? Describe how they are doing it.”. On the other hand, fine-grained questions focus on details such as the user’s rate of punch (Fig. 5 top) or whether the user’s back is rounded (Fig. 5 bottom). The feedbacks on these short-clips provide positive reinforcement, e.g., “This is looking great!” if the user is doing the exercise correctly (Fig. 5 top) and also provide useful advice, e.g., toe alignment while performing squats (Fig. 5 bottom).
Figure 5: Additional example annotations available on the short video clips from the Qevd-Fit-300K dataset (see also Figs. 2 and 1 in the main paper).
@@ Appendix C Additional Data Collection Details
Additional details pertaining to the various subsets of Qevd are provided in this section.
@@ C.1 Qevd-Fit-300K
Here we explore the labels associated with the short video clips in the Qevd-Fit-300K dataset. These labels are crucial for generating fitness questions and providing feedback for short-clip videos.
@@
Coarse and fine labels. As described in Sec. 3.2 in the main paper, the short video clips are annotated with coarse labels and fine-grained attributes in addition to the questions and feedbacks. Labels were verified for correctness by at least one additional human annotator. The full list of coarse labels and fine-grained attributes can be found at https://developer.qualcomm.com/software/ai-datasets/qevd.
@@ C.2 Qevd-Fit-Coach
Video Collection. Long-range workout videos were collected by letting users perform workout sessions consisting of three $1$-minute sections, where each section consisted of two exercises. Candidate exercises per section are shown in Tab. 5. Users were instructed to perform specific exercises for the indicated duration while facing the camera to imitate a virtual fitness coaching setup. They were also instructed to perform common mistakes and their corrections (similar to the short-clips). Temporally aligned textual coaching instructions were subsequently cleaned-up and verified using a simple web interface. All feedbacks were reviewed by at least one additional annotator.